In order to help identify baby growth patterns that are unusual, we need to construct a confidence interval estimate of the mean head circumference of all babies that are two months old. A random sample of 40 babies is obtained, and the mean head circumference is found to be 40.6 cm and a standard deviation is known to be 1.6 cm. Find a 99% confidence interval estimate of the mean head circumference of all two-month-old babies.

What is the margin of error?

Incorrect answer: 0.6515
Answer #1

Solution :

Given that,

Z/2 = 2.576

Margin of error = E = Z/2* ( /n)

= 2.576 * (1.6 / 40)

= 0.651639